# Discrete dark matter with light Dirac neutrinos

### v2

@@ Appendix C Neutrino mass matrix
$=$ $\begin{pmatrix}(m_{\nu})_{11}&(m_{\nu})_{12}&(m_{\nu})_{13}\\ (m_{\nu})_{21}&(m_{\nu})_{22}&(m_{\nu})_{23}\\ (m_{\nu})_{31}&(m_{\nu})_{32}&(m_{\nu})_{33}\end{pmatrix}$ (35)
@@
$(m_{\nu})_{11}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{1}+y_{1}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3}),$ (36)
$(m_{\nu})_{12}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{2}+y_{1}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3}),$ (37)
$(m_{\nu})_{13}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{3}+y_{1}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3}),$ (38)
$(m_{\nu})_{21}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{1}+y_{2}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3}),$ (39)
$(m_{\nu})_{22}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{2}+y_{2}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3}),$ (40)
$(m_{\nu})_{23}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{3}+y_{2}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3}),$ (41)
$(m_{\nu})_{31}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{1}+y_{3}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3}),$ (42)
$(m_{\nu})_{32}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{2}+y_{3}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3}),$ (43)

### v3

@@ Appendix C Neutrino mass matrix
$=$ $\begin{pmatrix}(m_{\nu})_{11}&(m_{\nu})_{12}&(m_{\nu})_{13}\\ (m_{\nu})_{21}&(m_{\nu})_{22}&(m_{\nu})_{23}\\ (m_{\nu})_{31}&(m_{\nu})_{32}&(m_{\nu})_{33}\end{pmatrix},$ (35)
@@
$(m_{\nu})_{11}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{1}+y_{1}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3});$ (36)
$(m_{\nu})_{12}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{2}+y_{1}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3});$ (37)
$(m_{\nu})_{13}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{1}y^{\prime}_{3}+y_{1}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3});$ (38)
$(m_{\nu})_{21}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{1}+y_{2}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\omega\mathcal{F}_{2}+\omega^{2}\mathcal{F}_{3});$ (39)
$(m_{\nu})_{22}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{2}+y_{2}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3});$ (40)
$(m_{\nu})_{23}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{2}y^{\prime}_{3}+y_{2}y^{\prime}_{3}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3});$ (41)
$(m_{\nu})_{31}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{1}+y_{3}y^{\prime}_{1}M_{N}(\mathcal{F}_{1}+\omega^{2}\mathcal{F}_{2}+\omega\mathcal{F}_{3});$ (42)
$(m_{\nu})_{32}$ $=$ $-\frac{v_{\eta}v_{\chi}}{2M_{N}}y_{3}y^{\prime}_{2}+y_{3}y^{\prime}_{2}M_{N}(\mathcal{F}_{1}+\mathcal{F}_{2}+\mathcal{F}_{3});$ (43)